OpenStudy (anonymous):

The sum of the fourth and twelfth term of an arithmetic progression is 20. What is the sum of the first 15 terms of the arithmetic progression?

OpenStudy (amistre64):

a4 + a12 = 20 that is 8 terms apart

OpenStudy (amistre64):

8(f+l)/2 = 20 8(f+l) = 40 f+l = 5

OpenStudy (amistre64):

im missing part there arent i ... or at least its a scaled version

OpenStudy (anonymous):

couldn't the sequence be 3,4,5,6,7,8,9,10,11,12,13,14,15,16,17? I'm sure there must be other answers, though.

OpenStudy (amistre64):

ninjas works

OpenStudy (anonymous):

like -4, -2, 0, 2, 4, 6, 8, 10, 12, 14, 16, 18, 20, 22, 24

OpenStudy (anonymous):

in both, the sum of the first 15 terms is 150.

OpenStudy (anonymous):

because if a4 + a12 is 20, a8 has to be 10, and a7 + a9 is 20, and so on. So the sum has to be 150.